OpenStudy (ineedhelpnow):

how do i solve x+9/5=x-10=11

OpenStudy (saifoo.khan):

the equation is incorrect.

OpenStudy (anonymous):

x+9/5=x-10 x-10=11 solve them seperately to get the answer

OpenStudy (anonymous):

is the question is correct than the answer will be same for both of them

OpenStudy (anonymous):

siffffff!!!!!!!!!!!!!!!!!!!!!!!!!!1

OpenStudy (ineedhelpnow):

oh! x+9/5=x-10/11 sorry.

OpenStudy (anonymous):

wah x canels out